OpenStudy (anonymous):

From a group of 5 freshmen, 3 sophomores, 4 juniors, and 3 seniors, how many 5-person committees are possible?

OpenStudy (anonymous):

5+3+4+3=15 total people, then 15*14*13*12*11 = 360360 possible arrangements on the basis that no person may hold more than one spot. However if repetitions are possible, there are a total of 15*15*15*15*15=759375 possible arrangements.